OpenStudy (anonymous):

Find the cube roots of 8(cos 216° + i sin 216°).

OpenStudy (anonymous):

you should use DeMoivre's theorem for complex numbers in their polar form

OpenStudy (anonymous):

\[r^3=8 \longrightarrow r=2\] \[3\theta = 216 + n360\] so solve this for n=0,1,2... to generate all the three roots